### v3

@@ 4.1 Medical Multimodal-Multitask Dataset Construction
Figure 2-(a) presents our general workflow for constructing the multimodal multitask medical datasets, which consists of four main steps: (1) medical tasks of interest definition; (2) task-specific multimodal data collection; (3) multimodal data processing and alignment; and (4) MQA construction. The details for the first two steps are described in Section 2.1. Here we introduce the third and fourth steps.
@@
The multimodal data processing to select qualified multimodal data and prepare them for the alignment, including the CT data processing, clinical data processing, and ground-truth label calculating. In CT data processing, we localize the sub-volumes that mainly contain the task-relevant regions in the 3D CT using a segmentation model [44]. Specifically, we segment three parts, i.e., left lung, right lung, and heart regions consisting of the myocardium, left/right atrium, left/right ventricle, and pulmonary artery. It is worth underlining that the precision of the segmentation results does not need to be extremely high. Our primary objective is to utilize rectangular boxes to wrap the segmented areas, ensuring that task-relevant sub-volumes are included and extraneous regions are disregarded. We excluded the CT series having less than 64 axial slices in all collected datasets. For each CT series, the reconstruction voxel sizes are prepared in the axial, coronal, and sagittal dimensions, which will be used as input to the CTViT. The clinical data processing is to represent various combinations of clinical data within free text. We have established a specific sentence format for each clinical element, as detailed in Table S1 of the supplementary. The final free-text clinical data for each examination is constructed by aggregating the sentences corresponding to all available and positive clinical data, as illustrated by the text inputs in Figures 1 and 5. In ground-truth label calculating, we first extract task-specific label information from different sources (see Table 1 for specific label sources of each task) and then combine all information to calculate the label. The details for each task ground-truth label calculation are described in Table S2 of the supplementary. Next, we align the clinical data in free text, the CT data with segmented parts and physical size, and the calculated labels for all exams in each task. In particular, each task anatomically corresponds to the segmented CT sub-volume. The key principle here is to remove the irrelevant image regions to reduce the computation cost while keeping the original resolution without losing the information. The task-specific CT sub-volumes are illustrated in the image input column of Table 1. Specifically, 2.5D left or right lung sub-volume is used as the image input for lung nodule detection and characterization. We fixed the number of slices to 16 for each input considering that a lung nodule is usually tiny relative to the whole lungs. For the nodule-presented case, the location labels of the left/right lung, slice number, and the bounding box coordinates are used to crop the target sub-volume. For the non-presented nodule case, the input sub-volume is randomly cropped within the segmented lung regions. The CVD tasks use a 3D rectangle box wrapping all heart regions as the image input. For lung cancer risk prediction, we separately input the left or right lung in a 3D rectangle box. In the training datasets, the location of the left and/or right lung where lung cancer is presented is required to align each lung with the risk labels. In the validation and test datasets, the location of lung cancer is not required as the ground-truth labels are patient-level and the predicted lung cancer risks are the maximum scores between two lungs for each patient. For all other chest disease diagnosis tasks, a 3D rectangle box wrapping both lungs is used as the image input. Subsequently, the MQA construction is to create questions and answers for each specific task with the aligned multimodal data, and the resulting MQA datasets define the model’s input and output formats. In Table 1, one example question and the corresponding answer candidates are presented for each task. In the training stage, ten different questions for each task were used as shown in Table S3 of the supplementary. Note that there is no need for extra labeling efforts by radiologists across the whole workflow so that large-scale medical multimodal multitask datasets can be systematically and cost-effectively constructed.
@@ 4.2.2 CTViT
CTViT extracts embedding features of multi-scale 3D CT volumes with physical size awareness. CTViT has two parts: a multi-scale CT tokenizer and an image encoder. To process a 3D CT scan, we divide each image volume into non-overlapped 3D patches as in [45]. Each 3D patch is referred to as an image token. Since various diseases are at different scales in the CT images, we design a multi-scale CT tokenizer, which consists of multiple linear embedding layers corresponding to different sizes of image patches, as shown in Figure 6-(a). Each embedding layer has a linear transformation and a set of learnable positional embeddings. Each image token embedding is the sum of its linear transformation and the positional embedding. All sizes of image tokens are mapped to the same image embedding space. Inspired by [41], we decompose the 3D position embedding into two parts indexing in-plane and through-plane positions respectively. In other words, we have two positional embeddings: one for the 2D space within each slice and the other for the 1D range of slice position. The 3D positional embedding is the sum of them. By doing so, the number of learned parameters can be reduced. Figure 6(c) shows the image encoder in detail. Different CT scans may have different physical sizes specific to the individual patient size. The physical size is an important factor in some clinical tasks. Thus, we encode the physical size for image tokens with sine-cosine functions of different frequencies and add it to the image token embedding. Then, the image encoder was implemented as the plain ViT [46] that consists of multiple self-attention Transformer layers and a subsequent linear transformation layer that maps the image embedding space to the task embedding space. By disentangling physical size from image contents, we can flexibly perceive any size of CT volumes with physical size awareness without resampling CT volumes to have a consistent image tensor across different inputs.
@@ 4.2.3 Text Transformer
Any decent language model can be utilized as the text Transformer in M3FM. Here the text Transformer was implemented as a Byte-level Byte-Pair-Encoding (BBPE) tokenizer [39], a text encoder consisting of the original Transformer layers [47], and a linear transformation layer, as shown in Figure 6-(d). On one hand, the text encoder encodes patient-specific clinical information, such as demographics, smoking history, disease history, cancer history, and other clinical data, which are free text; e.g., “The patient is 56.0 years old. Gender is Female. Ethnicity is neither Hispanic nor Latino. Height is 60.0 inches. Weight is 105.0 pounds. Education is associate degree/ some college. Former smoker. Smoking duration is 38.0 pack years. Smoking intensity is 20.0 cigarettes per day. 2.0 years since quit smoking. The patient had asthma (childhood) diagnosed at 7.0 years old. The patient had hypertension diagnosed at 53.0 years old. The patient had pneumonia diagnosed at 50.0 years old. Patient’s brother(s) (including half-brothers) have lung cancer.” On the other hand, the text encoder encodes free-text task instructions/questions, which are used as the input of the task encoder to extract task-specific embedding features from the multimodal data; e.g., “Is there any significant cardiovascular abnormality?” and “Predict the lung cancer risk over 6 years.” This approach allows for embedding any combination of clinical information through free-text prompting, regardless of order. The control signals for specific tasks are then extracted from the text prompts by the same text encoder. Again, the linear transform maps the text embedding space to the unified task embedding space. For all our M3FMs, the number of Transformer layers is 12, and the size of text token embeddings is 768.
@@ 4.2.4 Task Encoder
Figure 6-(e) illustrates the task encoder, designed to extract task-specific embedding features from the multimodal token embeddings, given the special token embedding. The task encoder was implemented with multiple Transformer layers and regards all tokens as a single input sequence. Note that only the special token is forwarded to the task encoder and the rest of the question tokens are ignored, as we empirically found that other tokens are not useful in practice. The special token embedding from the final Transformer layer serves as the task-specific embedding feature that integrates all multimodal data. For M3FM-Base, M3FM-Large, and M3FM-Huge, the number of Transformer layers is 4 in every case, and the sizes of task token embeddings are 768, 1,024, and 1,280, respectively.
@@ 4.4 Self-supervised Pretraining
A key step to optimize large models is self-supervised pretraining with large unlabeled datasets. In this study, we adapted the masked autoencoder method [40, 41] to pretrain our CTViT on the constructed OpenM3Chest pretraining dataset. Figure 6-(b) shows the CTViT pretraining architecture, which consists of CTViT, an image decoder, and a multi-scale linear de-tokenizer. The image encoder introduced in Subsection 4.2.2 was optimized by predicting masked cubes (85%) from a small number of visible cubes (15%). To reduce the memory overhead, only some selected slices along the longitude direction were predicted while recovering each 3D patch. We pre-trained CTViT with the defined multi-scale 3D CT volumes and a set of data augmentation operations, including random cropping, rotation, resizing, and perturbed display windowing. The text Transformer in M3FM was initialized with the off-the-shelf RoBERTa model pre-trained via masked language modeling and then trained end-to-end [49, 39].
@@ 4.6 Transfer Learning
M3FM is designed for adaptability and generalization, enabling the enhancement of out-of-distribution task performance through transfer learning. This capability extends to new tasks with varying image input dimensions, clinical data types, and output dimensions. To accommodate different image dimensions, the addition of a linear embedding layer suffices. For diverse clinical datasets, we can simply describe involved clinical data in free text to the model, without needing any modification on the M3FM architecture, as shown in Figure 5-(a). Specifically, adjusting to different output dimensions requires only the inclusion of a lightweight predictor. Consequently, M3FM can be easily fine-tuned to enable new tasks by leveraging the pre-trained model parameters.
@@ 4.8 Hardware Requirement
Data Availability The raw data of NLST are available from the NIH National Cancer Institute (https://cdas.cancer.gov/nlst/). The raw data of MIDRC are available at https://www.midrc.org/. Our curated OpenM3Chest dataset will be publicly available upon the acceptance of the paper. Restrictions apply to the availability of anonymized patient data from WFUSM and MGH for this project with institutional permission and are, thus, not publicly available.

### v4

@@ 4.1 Medical Multimodal-Multitask Dataset Construction
Figure 2 (a) presents our general workflow for constructing the multimodal multitask medical datasets, which consists of four main steps: (1) medical tasks of interest definition; (2) task-specific multimodal data collection; (3) multimodal data processing and alignment; and (4) MQA construction. The details for the first two steps are described in Section 2.1. Here we introduce the third and fourth steps.
@@
The multimodal data processing is to select qualified multimodal data and prepare them for the alignment, including the CT data processing, clinical data processing, and ground-truth label calculating. In CT data processing, we localize the sub-volumes that mainly contain the task-relevant regions in the 3D CT using a segmentation model [46]. Specifically, we segment three parts, i.e., left lung, right lung, and heart regions consisting of the myocardium, left/right atrium, left/right ventricle, and pulmonary artery. It is worth underlining that the precision of the segmentation results does not need to be extremely high. Our primary objective is to utilize rectangular boxes to wrap the segmented areas, ensuring that task-relevant sub-volumes are included and extraneous regions are disregarded. We excluded the CT series having less than 64 axial slices in all collected datasets. For each CT series, the reconstruction voxel sizes are prepared in the axial, coronal, and sagittal dimensions, which will be used as input to the CTViT. The clinical data processing is to represent various combinations of clinical data within free text. We have established a specific sentence format for each clinical element, as detailed in Table S1 of the supplementary. The final free-text clinical data for each examination is constructed by aggregating the sentences corresponding to all available and positive clinical data, as illustrated by the text inputs in Figures 1 and 7. In ground-truth label calculating, we first extract task-specific label information from different sources (see Table 1 for specific label sources of each task) and then combine all information to calculate the label. The details for each task ground-truth label calculation are described in Table S2 of the supplementary. Next, we align the clinical data in free text, the CT data with segmented parts and physical size, and the calculated labels for all exams in each task. In particular, each task anatomically corresponds to the segmented CT sub-volume. The key principle here is to remove the irrelevant image regions to reduce the computation cost while keeping the original resolution without losing the information. The task-specific CT sub-volumes are illustrated in the image input column of Table 1. Specifically, 2.5D left or right lung sub-volume is used as the image input for lung nodule detection and characterization. We fixed the number of slices to 16 for each input considering that a lung nodule is usually tiny relative to the whole lungs. For the nodule-presented case, the location labels of the left/right lung, slice number, and the bounding box coordinates are used to crop the target sub-volume. For the non-presented nodule case, the input sub-volume is randomly cropped within the segmented lung regions. The CVD tasks use a 3D rectangle box wrapping all heart regions as the image input. For lung cancer risk prediction, we separately input the left or right lung in a 3D rectangle box. In the training datasets, the location of the left and/or right lung where lung cancer is presented is required to align each lung with the risk labels. In the validation and test datasets, the location of lung cancer is not required as the ground-truth labels are patient-level and the predicted lung cancer risks are the maximum scores between two lungs for each patient. For all other chest disease diagnosis tasks, a 3D rectangle box wrapping both lungs is used as the image input. Subsequently, the MQA construction is to create questions and answers for each specific task with the aligned multimodal data, and the resulting MQA datasets define the model’s input and output formats. In Table 1, one example question and the corresponding answer candidates are presented for each task. In the training stage, ten different questions for each task were used as shown in Table S3 of the supplementary. Note that there is no need for extra labeling efforts by radiologists across the whole workflow so that large-scale medical multimodal multitask datasets can be systematically and cost-effectively constructed.
@@ 4.2.2 CTViT
CTViT extracts embedding features of multi-scale 3D CT volumes with physical size awareness. CTViT has two parts: a multi-scale CT tokenizer and an image encoder. To process a 3D CT scan, we divide each image volume into non-overlapped 3D patches as in [47]. Each 3D patch is referred to as an image token. Since various diseases are at different scales in the CT images, we design a multi-scale CT tokenizer, which consists of multiple linear embedding layers corresponding to different sizes of image patches, as shown in Figure 8 (a). Each embedding layer has a linear transformation and a set of learnable positional embeddings. Each image token embedding is the sum of its linear transformation and the positional embedding. All sizes of image tokens are mapped to the same image embedding space. Inspired by [42], we decompose the 3D position embedding into two parts indexing in-plane and through-plane positions respectively. In other words, we have two positional embeddings: one for the 2D space within each slice and the other for the 1D range of slice position. The 3D positional embedding is the sum of them. By doing so, the number of learned parameters can be reduced. Figure 8(c) shows the image encoder in detail. Different CT scans may have different physical sizes specific to the individual patient size. The physical size is an important factor in some clinical tasks. Thus, we encode the physical size for image tokens with sine-cosine functions of different frequencies and add it to the image token embedding. Then, the image encoder was implemented as the plain ViT [48] that consists of multiple self-attention Transformer layers and a subsequent linear transformation layer that maps the image embedding space to the task embedding space. By disentangling physical size from image contents, we can flexibly perceive any size of CT volumes with physical size awareness without resampling CT volumes to have a consistent image tensor across different inputs.
@@ 4.2.3 Text Transformer
Any decent language model can be utilized as the text Transformer in M3FM. Here the text Transformer was implemented as a Byte-level Byte-Pair-Encoding (BBPE) tokenizer [40], a text encoder consisting of the original Transformer layers [49], and a linear transformation layer, as shown in Figure 8 (d). On one hand, the text encoder encodes patient-specific clinical information, such as demographics, smoking history, disease history, cancer history, and other clinical data, which are free text; e.g., “The patient is 56.0 years old. Gender is Female. Ethnicity is neither Hispanic nor Latino. Height is 60.0 inches. Weight is 105.0 pounds. Education is associate degree/ some college. Former smoker. Smoking duration is 38.0 pack years. Smoking intensity is 20.0 cigarettes per day. 2.0 years since quit smoking. The patient had asthma (childhood) diagnosed at 7.0 years old. The patient had hypertension diagnosed at 53.0 years old. The patient had pneumonia diagnosed at 50.0 years old. Patient’s brother(s) (including half-brothers) have lung cancer.” On the other hand, the text encoder encodes free-text task instructions/questions, which are used as the input of the task encoder to extract task-specific embedding features from the multimodal data; e.g., “Is there any significant cardiovascular abnormality?” and “Predict the lung cancer risk over 6 years.” This approach allows for embedding any combination of clinical information through free-text prompting, regardless of order. The control signals for specific tasks are then extracted from the text prompts by the same text encoder. Again, the linear transform maps the text embedding space to the unified task embedding space. For all our M3FMs, the number of Transformer layers is 12, and the size of text token embeddings is 768.
@@ 4.2.4 Task Encoder
Figure 8 (e) illustrates the task encoder, designed to extract task-specific embedding features from the multimodal token embeddings, given the special token embedding. The task encoder was implemented with multiple Transformer layers and regards all tokens as a single input sequence. Note that only the special token is forwarded to the task encoder and the rest of the question tokens are ignored, as we empirically found that other tokens are not useful in practice. The special token embedding from the final Transformer layer serves as the task-specific embedding feature that integrates all multimodal data. For M3FM-Base, M3FM-Large, and M3FM-Huge, the number of Transformer layers is 4 in every case, and the sizes of task token embeddings are 768, 1,024, and 1,280, respectively.
@@ 4.4 Self-supervised Pretraining
A key step to optimize large models is self-supervised pretraining with large unlabeled datasets. In this study, we adapted the masked autoencoder method [41, 42] to pretrain our CTViT on the constructed OpenM3Chest pretraining dataset. Figure 8 (b) shows the CTViT pretraining architecture, which consists of CTViT, an image decoder, and a multi-scale linear de-tokenizer. The image encoder introduced in Subsection 4.2.2 was optimized by predicting masked cubes (85%) from a small number of visible cubes (15%). To reduce the memory overhead, only some selected slices along the longitude direction were predicted while recovering each 3D patch. We pre-trained CTViT with the defined multi-scale 3D CT volumes and a set of data augmentation operations, including random cropping, rotation, resizing, and perturbed display windowing. The text Transformer in M3FM was initialized with the off-the-shelf RoBERTa model pre-trained via masked language modeling and then trained end-to-end [51, 40].
@@ 4.6 Transfer Learning
M3FM is designed for adaptability and generalization, enabling the enhancement of out-of-distribution task performance through transfer learning. This capability extends to new tasks with varying image input dimensions, clinical data types, and output dimensions. To accommodate different image dimensions, the addition of a linear embedding layer suffices. For diverse clinical datasets, we can simply describe involved clinical data in free text to the model, without needing any modification on the M3FM architecture, as shown in Figure 7 (a). Specifically, adjusting to different output dimensions requires only the inclusion of a lightweight predictor. Consequently, M3FM can be easily fine-tuned to enable new tasks by leveraging the pre-trained model parameters.
@@ 4.8 Hardware Requirement
Data Availability The raw data of NLST are available in the NIH/NCI Cancer Data Access System (https://cdas.cancer.gov/nlst/). The raw data of MIDRC are available at the MIDRC’s Data Access Portal (https://www.midrc.org/). Our curated OpenM3Chest dataset will be made publicly available upon the acceptance of the paper. Restrictions apply to the availability of the anonymized patient data from WFUSM and MGH for this project with institutional permission, while these data could be shared with potential users through negotiation with each of the institutes.